# ELTEX: A Framework for Domain-Driven Synthetic Data Generation

### v1

@@ Abstract
We present ELTEX (Efficient LLM Token Extraction), a domain-driven framework for generating high-quality synthetic training data in specialized domains. While Large Language Models (LLMs) have shown impressive general capabilities, their performance in specialized domains like cybersecurity remains limited by the scarcity of domain-specific training data. ELTEX addresses this challenge by systematically integrating explicit domain indicator extraction with dynamic prompting to preserve critical domain knowledge throughout the generation process. We demonstrate ELTEX’s effectiveness in the context of blockchain-related cyberattack detection, where we fine-tune Gemma-2B using various combinations of real and ELTEX-generated data. Our results show that the ELTEX-enhanced model achieves performance competitive with GPT-4 across both standard classification metrics and uncertainty calibration, while requiring significantly fewer computational resources. We release a curated synthetic dataset of social media texts for cyberattack detection in blockchain. Our work demonstrates that domain-driven synthetic data generation can effectively bridge the performance gap between resource-efficient models and larger architectures in specialized domains.
@@ 1 Introduction
Large Language Models (LLMs) have achieved state-of-the-art performance across various tasks, extending beyond traditional language processing (MMLU-Pro Wang et al. (2024), LiveCodeBench; Jain et al. (2024)), yet adapting them to specialized domains remains challenging Ling et al. (2024). This limitation is particularly evident in cybersecurity, where benchmarks such as CyberBench Liu et al. (2024) reveal that even state-of-the-art LLMs (e.g., GPT-4) struggle with domain-specific tasks, including cyber-threat named entity recognition and risk mitigation planning.
Early signals of emerging cyberattacks frequently appear in unstructured natural language channels: social media platforms, dark web forums, and technical communities, long before they are formally documented in vulnerability databases Paladini et al. (2024); Almukaynizi et al. (2017); Khandpur et al. (2017). This makes NLP crucial in cyber threat intelligence, as it enables the collection and analysis of evidence-based indicators of compromise to identify and mitigate emerging attacks Arazzi et al. (2023); Sufi (2023).
@@
While social media platforms provide abundant cybersecurity-related data, recent studies highlight significant challenges in data quality, including duplicate content Mu et al. (2024), inconsistent labeling Zong et al. (2019), and limited coverage of diverse attack types Almahmoud et al. (2023). In response, LLM-based synthetic data generation, has emerged as a promising solution Li et al. (2023), offering potential for overcoming data quality issues while preserving privacy and domain validity. Existing approaches to synthetic data generation have primarily relied on implicit knowledge transfer through few-shot examples or post-generation validation Patil and Gudivada (2024). Recent advances have introduced more sophisticated methodologies: UniGen Wu et al. (2024) implements attribute-guided generation with RAG-based validation, while CRAFT Ziegler et al. (2024) employs corpus-based retrieval with few-shot learning. Despite these advances, these methods still face challenges in extracting and preserving domain-specific knowledge.
We address this gap with Efficient LLM Token Extraction (ELTEX), a domain-driven framework that systematically integrates explicit domain indicator extraction with dynamic prompting. ELTEX preserves critical domain nuances throughout the generation process. While we demonstrate ELTEX’s effectiveness through text classification of blockchain-related cyberattack discussions on social media, the framework’s methodology is potentially generalizable to other specialized domains.
To validate ELTEX, we fine-tune Gemma-2B Team et al. (2024b) using various combinations of real and ELTEX-generated data. We demonstrate that our domain-driven approach enables this lightweight model to achieve performance competitive with GPT-4o OpenAI (2023) across both standard classification metrics and uncertainty calibration measured by the Brier score Brier (1950), while requiring significantly fewer computational resources.
Our main contributions are as follows:
  1. 1.
    Domain-Driven Synthetic Data Generation. We propose ELTEX, a framework that systematically integrates explicit domain indicator extraction with dynamic prompt refinement to generate high-fidelity, context-grounded synthetic data.
  2. 2.
    Downstream Evaluation. Through fine-tuning experiments on Gemma-2B, we demonstrate that domain-driven synthetic data can bridge the performance gap between resource-efficient models and larger architectures in specialized domains.
  3. 3.
    Released Synthetic Dataset. We release a carefully curated synthetic dataset of 11,448 social media texts for cyberattack detection in blockchain.  [^1] The dataset is available on HuggingFace
@@ 2 Related Work
2.1 Domain-Adapted Small Models
While large-scale models (e.g., GPT-4) offer strong general performance, recent benchmarks such as CyberBench Liu et al. (2024) reveal that fine-tuned, smaller models can rival or even surpass their larger counterparts in cybersecurity-specific tasks. For instance, Llama-2-13B fine-tuned with CyberInstruct surpasses GPT-4 in certain metrics (e.g., 72.6 vs. 63.0 for common vulnerabilities and exposures accuracy).
These findings demonstrate that domain-adapted smaller models can match larger ones, motivating our choice to explore more compact models’ capabilities through Gemma-2B. Smaller domain-adapted models offer practical advantages for deployment, particularly when enhanced with techniques like LoRA for domain adaptation. This approach addresses key requirements in cybersecurity applications where privacy constraints and real-time monitoring demands necessitate efficient, local inference capabilities.
2.2 LLMs for Cyber Threat Intelligence
LLMs have recently attracted attention for automated threat intelligence, vulnerability detection, and anomaly analysis Perrina et al. (2023); Cheshkov et al. (2023); Karlsen et al. (2024). Despite their potential, three interrelated obstacles hinder their deployment in cybersecurity:
  1. 1.
    Rapid Domain Drift: Emerging cyber threats can render previously trained models obsolete Cremer et al. (2022). Attack vectors evolve faster than in many other NLP settings, forcing frequent re-training or advanced domain-adaptation strategies to maintain relevance.
  2. 2.
    Data Volume and Velocity: The sheer scale of social media demands computationally efficient real-time analysis Zhou et al. (2024). High-volume API-based LLM inference can lead to unsustainable costs and latency issues.
  3. 3.
    Privacy and Proprietary Constraints: Third-party LLM APIs pose privacy risks Yan et al. (2024), making local or on-premise solutions appealing for many organizations. However, smaller local models may initially underperform without domain-specific adaptation.
@@ 2.2 LLMs for Cyber Threat Intelligence
Recent research tackles these barriers through two complementary approaches: (i) high-fidelity synthetic data generation to alleviate domain-specific data scarcity and keep pace with emerging threats Ferrag et al. (2024), and (ii) compact model architectures (e.g., Microsoft’s phi-3-mini Abdin et al. (2024)), which can be deployed locally to reduce inference costs, preserve privacy, and facilitate real-time monitoring.
2.3 LLM-based Synthetic Data Generation for Domain-Specific Tasks
@@
The effectiveness of domain adaptation and small models heavily depends on the quality and availability of training data. Social media sources complicate this need: duplicated content, inconsistent terminology, and topical imbalances frequently undermine data reliability Mu et al. (2024); Zong et al. (2019); Almahmoud et al. (2023). Coupled with the rapid evolution of cyber threats, manual dataset curation and continuous updates become prohibitively difficult. LLM-based synthetic data generation offers a promising workaround by producing privacy-preserving samples rich in domain-specific details. Conventional approaches rely on heuristic prompts that combine task specifications, generation conditions, and in-context examples to drive LLM output Long et al. (2024). For more intricate datasets, multi-step generation techniques decompose the process into manageable subtasks.
@@
Frameworks, such as UniGen Wu et al. (2024) and CRAFT Ziegler et al. (2024), have aimed to streamline synthetic data generation, though each faces distinct challenges in knowledge-intensive domains. UniGen provides a unified approach to generating diverse datasets. CRAFT leverages corpus retrieval and LLM augmentation to synthesize task-specific data from few-shot examples, demonstrating success in QA and summarization tasks. However, these frameworks face limitations in knowledge-intensive domains: UniGen struggles with specialized knowledge retention, and CRAFT’s reliance on public web corpora limits its efficacy for cybersecurity where domain-specific data is scarce and sensitive. Our proposed ELTEX framework builds on these foundations and bridges the gap by systematically extracting and integrating explicit domain knowledge, particularly benefiting knowledge-intensive applications where existing approaches fall short.
@@ 3 ELTEX Framework
ELTEX, as illustrated in Figure 1, consists of five main components: (1) sample data collection and deduplication, (2) token extraction prompt construction, (3) synthetic data generation, (4) final deduplication and (5) post-generation quality assurance (QA).
Figure 1: Efficient LLM Token Extraction Pipeline
@@ 3.1 Data Collection
Data Sources and Collection
Our research focused on creating a dataset specifically for the classification of social media messages related to cyberattacks in the blockchain industry. By utilizing the official API (v2) provided by X (formerly known as Twitter) X (2025), we obtained data on cyberattacks and general topics from the past three years within the blockchain ecosystem. We curated a list of documented cyberattacks on various blockchain participants to systematically target relevant events. For each entry in this list, we defined a search window around the reported start date to capture discussions that potentially reflect early indicators or live commentary on the attack (see  4.1 for attack types distribution).
Overall, we retrieved 1,766 real data samples (see Table 1 for detailed statistics throughout the pipeline). This initial dataset captures both high-signal messages linked to specific attacks and broader blockchain discussions that do not pertain to cyberattacks.
All data collection procedures were conducted according to ethical guidelines (see Section Ethical Considerations).
Annotation Process
Collected messages underwent a two-stage annotation process:
  1. 1.
    Preliminary Annotation: GPT-4o (version August 6, 2024) generated classification scores ranging from 0 to 1 using the predefined prompt.
  2. 2.
    Human Review: In-house annotators refined labels, guided by model-provided scores.
The full annotation process, including the prompt and validation methods, is detailed in Appendix B.
Deduplication Process
To ensure data diversity, annotated messages were deduplicated using a similarity threshold of 0.9 (See Section 3.6). This threshold was selected based on experimental results (detailed in Appendix D) to balance diversity preservation with duplicate removal. The process reduced cyberattack messages by 11.8% and general messages by 5.2% (see Table 1).
@@ 3.2 Token Extraction Prompt Construction
To effectively guide GPT-4o (version August 6, 2024) in generating domain-relevant tokens for early cyberattack detection, we designed a specialized extraction prompt. The prompt integrates instructions on the expected output format, along with curated lists of potential cyberattack indicators and concise examples of social media messages. Our approach is inspired by recent work on training data extraction from language models Yu et al. (2023), which demonstrates that carefully engineered prompts can coax latent training information from a model. Building on these insights, we focused on extracting key domain-specific tokens for cyberattacks in blockchain.
Indicator Generation and Validation
The process of generating and validating indicators consists of the following steps:
  1. 1.
    Initial Indicator Compilation: Query multiple LLMs (e.g., GPT-4o, Claude 3.5 Anthropic (2023), Gemini 1.5 Team et al. (2024a)) to generate a broad list of candidate indicators, such as suspicious transaction volume spikes or compromised private keys. Using multiple models ensures comprehensive coverage, as each model may introduce unique terminology or previously overlooked signals.
  2. 2.
    Summarization and Refinement: Iteratively summarize and consolidate the initial indicator list to merge semantically similar items and remove redundancies. This process leverages LLMs to distill the most relevant indicators while reducing noise, minimizing the need for extensive human expert review.
  3. 3.
    Stabilization and Validation: Perform two to three summarization iterations until the indicator list stabilizes. The final set is then reviewed for relevance and accuracy, ensuring alignment with domain-specific requirements.
For more detailed insights into this process, refer to Appendix A.
Final Prompt Construction
We structured the final prompt to include:
  1. 1.
    Task Description & Critical Instructions: Outlining the goal of generating early warning messages for potential cyberattacks and specifying constraints (e.g., anonymizing non-government entities).
  2. 2.
    Indicators List: A condensed set of validated attack indicators, guiding the model’s attention to relevant signals.
  3. 3.
    In-Context Examples: Real, anonymized social media messages.
By leveraging these elements within the ELTEX pipeline, we reduce reliance on time-consuming expert consultations and systematically incorporate both broad linguistic knowledge from LLMs and domain-specific insights. A detailed example of the constructed prompt, along with a broader discussion of indicator generation and multi-LLM fusion, is provided in Appendix A.
@@ 3.3 Synthetic Data Generation
We employed the GPT-4o (version August 6, 2024) model via the Azure OpenAI API (version 2024-05-01-preview) Microsoft (2025) to generate domain-relevant synthetic data for early cyberattack detection. While GPT-4o was selected for its status as a flagship LLM at the time, ELTEX is designed to be compatible with any large-scale language model that offers high-quality generation and a recent knowledge cut-off date, allowing users to choose the most suitable provider and model based on their specific needs.
Batching and Prompting
The generation process began by shuffling and batching real data samples into groups of 10 messages. Each batch was paired with a predefined ELTEX prompt, and for every API request, the model was configured to generate 100 synthetic messages. Due to the stochastic nature of LLM generation, the actual number of messages per request could slightly vary. To ensure efficient downstream processing, we utilized the API’s structured output feature OpenAI (n.d.), receiving the synthetic results in a JSON array. This process resulted in approximately 16,030 initial synthetic samples.
Temperature Settings
The model’s generation temperature was set to 0.8. Although varying the temperature between 0.0 and 1.0 resulted in only moderate changes in the overall diversity of generated messages, the deduplication process revealed important practical implications. At lower temperatures (0.0–0.6), applying a 0.9 similarity threshold reduced the final dataset by 50% or more, risking the loss of important thematic signals. In contrast, higher temperatures (0.9–1.0) often produced outputs that deviated from the desired format. Therefore, a temperature of 0.8 was selected to retain a sufficiently large and diverse dataset. Detailed results, including retention rates at different temperature settings, are provided in Appendix D.
3.4 Deduplication Process
All newly generated messages underwent a deduplication step based on embedding similarity before storage in a database (See Section 3.6). We adopted a similarity threshold of 0.9, with messages exceeding this threshold considered duplicates or near-duplicates. Experiments exploring alternative thresholds (e.g., 0.8) are detailed in Appendix D. While lower thresholds increase lexical and semantic diversity, they may inadvertently remove valuable linguistic nuances. Striking an optimal balance between diversity and utility remains a subject for future research. After applying the 0.9 threshold, we retained 11,465 messages (approximately 71.5% of the initially generated data).
@@ 3.5 Post-generation QA
We performed an additional QA step using NotebookLM Google (2025) to eliminate misclassified messages. NotebookLM was chosen for its advanced RAG capabilities, which effectively filter out irrelevant messages. Specifically, we uploaded the generated synthetic dataset to NotebookLM and instructed the LLM to identify and flag any contextually irrelevant content to its respective topic. The flagged messages were iteratively removed until no further content was detected that deviated from the predefined criteria. This approach can be adapted to any domain by specifying custom irrelevant text patterns. The final synthetic dataset contained 11,448 entries, as shown in Table 1.
@@
A cost analysis of the GPT-4o-based synthetic data generation pipeline, including token usage, is provided in Appendix F.
Clustering Analysis
We performed semantic clustering using the $k$-means algorithm MacQueen (1967) on the embeddings of cyberattack-related messages. The optimal number of clusters ($k=10$) was determined through silhouette score optimization, although the smooth elbow curve suggested no distinct natural clustering tendency in the embedding space. This indicates that the selection of $k$ should be viewed as an analytical convenience rather than a reflection of inherent semantic boundaries. We visualized the resulting clusters using UMAP.
In the real dataset, clusters were distinct and often aligned with specific incidents (e.g., Euler Finance exploit, Poloniex hack). Prominent keywords such as “million,” “hack,” and platform-specific terms suggested strong event alignment, though some overlap reflected the semantically rich nature of real data. In the synthetic dataset, clusters were more diffuse, with overlap and elongated structures, indicating broad but less incident-specific coverage. Keyword analysis highlighted additional abstract indicators — such as “transaction,” “breach,” “alert,” and “phishing” — not common in real data. While these broader terms captured potential cyberattacks, they lacked the event-specific nuance of the real dataset.
These findings suggest real data preserves detailed, incident-specific features, whereas synthetic data captures a wider array of abstract indicators. A hybrid dataset could combine these strengths, motivating our fine-tuning of Gemma-2B on combined real and synthetic samples for enhanced cybersecurity domain adaptation. Complete clustering results and visualizations are provided in Appendix D.
3.6 Deduplication Methodology
Our deduplication pipeline (details in Appendix E) employs a two-stage process to ensure data quality and diversity. First, exact-match filtering removes verbatim duplicates within each batch. Next, each message is embedded using the BGE-base-en-v1.5 model Xiao et al. (2023) and compared against existing embeddings via cosine similarity. Messages are retained only if their highest similarity score falls below a set threshold (e.g., 0.9). This strategy effectively filters near-duplicates while preserving meaningful variations without incurring excessive computational costs.
4 Experimental Setup
4.1 Dataset Construction
Our training pipeline leverages the dataset of 1,603 real-world social media messages, which we first split into training (80%) and validation (20%) sets. Using ELTEX, we expanded only the training portion into a synthetic dataset of 8,892 messages (5,530 cyberattack, 3,362 general), with generated classification scores ranging from 0 to 1. We maintained the original classdistribution while introducing diverse attack scenarios. This approach ensures no data leakage between the validation set and synthetic data generation. r model’s generalization capabilities, we curated training and and test datasets that maintain temporal separation and cover diverse attack vectors. In total, we collected an independent test set of 398 real-world messages (206 labeled as cyberattack and 192 as general), drawn from entirely distinct attack events not present in our training data. Table 2 shows the distribution of primary attack types across both training and test sets, illustrating our focus on a broad range of scenarios (e.g., DeFi exploits, wallet breaches, social engineering) to ensure robust evaluation

### v2

@@ Abstract
We introduce Efficient LLM Token Extraction (ELTEX), a framework addressing the critical challenge of LLM domain specialization by systematically extracting and integrating domain indicators throughout synthetic data generation. Unlike approaches relying on implicit knowledge transfer, ELTEX explicitly leverages domain signals to maintain specialized knowledge integrity. In our cybersecurity case study, ELTEX-enhanced data enables a fine-tuned Gemma-2B model to achieve performance competitive with GPT-4o on blockchain cyberattack classification while reducing computational requirements. Our Google Sheets implementation makes ELTEX accessible to non-technical users. Our contributions include: (1) the ELTEX framework; (2) Google Sheets Add-on implementation; (3) empirical validation showing how ELTEX bridges performance gaps between small and large models; and (4) a synthetic dataset of 11,448 texts for blockchain cyberattack detection.
@@ 1 Introduction
Domain specialization remains a critical challenge for deploying Large Language Models (LLMs) in industry settings Ling et al. (2024). While LLMs demonstrate impressive general capabilities, their performance deteriorates when confronted with specialized terminology and contextual nuances found in professional domains Kim et al. (2024); Lu et al. (2024).
We introduce Efficient LLM Token Extraction (ELTEX), a domain-driven framework that systematically integrates explicit domain indicator extraction with dynamic prompting to preserve critical domain nuances throughout the synthetic data generation process. ELTEX addresses a fundamental limitation in existing approaches: the difficulty of maintaining domain-specific knowledge integrity when generating training data for specialized applications. Unlike general-purpose data generation, ELTEX explicitly identifies and leverages domain indicators to guide the generation process.
@@
Existing approaches to synthetic data generation have primarily relied on implicit knowledge transfer through few-shot examples or post-generation validation Patil and Gudivada (2024). Recent advances have introduced more sophisticated methodologies: UniGen Wu et al. (2024) implements attribute-guided generation with RAG-based validation, while CRAFT Ziegler et al. (2024) employs corpus-based retrieval with few-shot learning. However, these methods still face challenges in explicitly extracting and preserving domain-specific knowledge throughout the generation pipeline.
The effectiveness of ELTEX is particularly evident in cybersecurity applications, where benchmarks such as CyberBench Liu et al. (2024) reveal that even state-of-the-art LLMs (e.g., GPT-4) struggle with domain-specific tasks. For example, in our case study, we demonstrate ELTEX’s effectiveness through text classification of blockchain-related cyberattack discussions, enabling a fine-tuned Gemma-2B model Team et al. (2024b) to achieve performance competitive with GPT-4o while requiring significantly fewer computational resources.
To enhance accessibility, we have deployed ELTEX as a Google Sheets add-on, making synthetic data generation accessible to non-technical users. This deployment has significantly reduced the time required to create specialized training datasets, enabling rapid response to emerging domain-specific patterns across various industries.
Our contributions include: (i) ELTEX, a framework that systematically identifies and leverages domain indicators to generate high-fidelity, context-grounded synthetic data; (ii) a Google Sheets add-on enabling non-technical users to generate and manage datasets;  [^1] ELTEX Add-on (iii) empirical validation showing that ELTEX-generated data bridges the performance gap between small and large models in specialized tasks; and (iv) a carefully curated synthetic dataset of 11,448 social media texts for cyberattack detection in blockchain.  [^2] The dataset is available on HuggingFace
@@ 2 Related Work
Specialized domains requiring LLM adaptation often share three critical barriers: (i) privacy constraints that limit data sharing and collection Subramani et al. (2023), (ii) rapid domain drift as terminology, threats, and regulations continuously evolve Lu et al. (2018), and thus (iii) insufficient labeled data volume to capture the complexity and nuance of domain-specific tasks. These obstacles create a significant gap between LLMs’ general capabilities and their effectiveness in specialized contexts.
@@
Recent research tackles these barriers through two complementary approaches: (i) high-fidelity synthetic data generation to alleviate domain-specific data scarcity Ferrag et al. (2024), and (ii) compact model architectures (e.g., Microsoft’s phi-3-mini Abdin et al. (2024)), which can be deployed locally to reduce inference costs, preserve privacy and increase throughput.
@@
LLM-based synthetic data generation offers a promising workaround by producing privacy-preserving samples rich in domain-specific details. Conventional approaches rely on heuristic prompts that combine task specifications, generation conditions, and in-context examples to drive LLM output Long et al. (2024). For more intricate datasets, multi-step generation techniques decompose the process into manageable subtasks.
@@
Frameworks, such as UniGen Wu et al. (2024) and CRAFT Ziegler et al. (2024), have aimed to streamline synthetic data generation, though each faces distinct challenges in knowledge-intensive domains. UniGen provides a unified approach to generating diverse datasets. CRAFT leverages corpus retrieval and LLM augmentation to synthesize task-specific data from few-shot examples, demonstrating success in QA and summarization tasks. However, these frameworks face limitations in knowledge-intensive domains: UniGen struggles with specialized knowledge retention, and CRAFT’s reliance on public web corpora limits its efficacy where domain-specific data is scarce and sensitive. Our proposed ELTEX framework builds on these foundations and bridges the gap by systematically extracting and integrating explicit domain knowledge, particularly benefiting knowledge-intensive applications where existing approaches fall short.
@@ 3 ELTEX Framework
ELTEX is designed to address domain specialization challenges across industries. As a case study, we demonstrate its application to cybersecurity for detecting blockchain-related cyberattacks in social media. The framework consists of five components (Figure 1): (1) data collection, (2) token extraction prompt construction, (3) synthetic data generation, (4) deduplication, and (5) post-generation quality assurance (QA). Below, we detail each step using cybersecurity as an illustrative example while emphasizing design choices that ensure adaptability to other domains.
Figure 1: ELTEX Pipeline
@@ 3.1 Data Collection
ELTEX requires seed data to guide generation, with collection methods adaptable to industry context. Organizations can leverage: (1) curated seed examples, (2) or synthetic bootstrapping, when even seed examples are scarce, ELTEX can bootstrap from descriptive prompts alone.
For the case study, we used the official X (formerly Twitter) API (v2) X (2025) to collect data on cyberattacks from the past three years within the blockchain ecosystem. We retrieved 1,766 initial samples (see Table 1 for statistics). These messages underwent a two-stage annotation process: (1) preliminary annotation using GPT-4o and (2) human review to refine labels. The full annotation process details are in Appendix B. All data collection procedures followed ethical guidelines (see Section Ethical Considerations).
For deduplication, we embedded each message using BGE-base-en-v1.5 Xiao et al. (2023) and removed samples exceeding a 0.9 vector similarity threshold. We selected this threshold based on pilot experiments indicating that lower values (e.g., 0.8) eliminated too many valid variations (detailed in Appendix D).
@@ 3.2 Token Extraction Prompt Construction
The core innovation of ELTEX is its systematic approach to domain knowledge preservation through explicit token extraction. Unlike approaches relying solely on implicit knowledge transfer via few-shot examples, ELTEX actively identifies and incorporates domain-specific indicators into the generation process. The token extraction process involves querying multiple LLMs to generate candidate domain indicators, followed by consolidation to merge semantically similar items and remove redundancies. This multi-model approach ensures comprehensive coverage, as each LLM may introduce overlooked signals.
We designed a specialized extraction prompt to effectively guide GPT-4o in generating domain-relevant tokens for early cyberattack detection. The prompt integrates instructions on the expected output format, curated lists of potential cyberattack indicators and examples of social media messages.
For the case study, we used GPT-4o, Claude 3.5 Anthropic (2023), and Gemini 1.5 Team et al. (2024a) to generate candidate indicators (e.g., unusual spikes in transaction volume, abnormal confirmation times, unexpected changes in mining difficulty). A detailed example of the constructed prompt, along with a broader discussion of indicator generation and multi-LLM fusion, is provided in Appendix A.
@@ 3.3 Synthetic Data Generation
This process is designed to be model-agnostic and compatible with any LLM offering high-quality generation capabilities. The generation approach includes shuffling and batching seed data samples, pairing each batch with the ELTEX prompt containing extracted domain tokens.
Temperature settings moderately influence the diversity and format of synthetic data. Lower temperatures (0.0–0.6) can yield repetitive outputs, medium temperatures (0.7–0.8) balance diversity with domain consistency, and high temperatures (0.9–1.0) often produce outputs that deviate from the desired format. Therefore, we selected a temperature of 0.8 to retain a sufficiently large and diverse dataset. Detailed results, including retention rates at different temperature settings, are provided in Appendix D.
For the case study, we employed GPT-4o via the Azure OpenAI API Microsoft (2025). We shuffled and batched real data samples into groups of 10 messages, paired each batch with the ELTEX prompt, and configured the model to generate 100 synthetic messages per request with a temperature of 0.8. To ensure efficient downstream processing, we utilized structured output formatting OpenAI (n.d.), receiving results in a JSON array. Due to the stochastic nature of LLM generation, the actual number of messages per request may vary. This process yielded approximately 16,030 initial synthetic samples. All newly generated messages underwent the same deduplication step described in 3.1. A cost analysis of the GPT-4o-based synthetic data generation pipeline, including token usage, is provided in Appendix F.
@@ 3.4 Post-generation QA
While our Google Sheets add-on lacks an integrated QA module, we used NotebookLM Google (2025) for this specific case study as a non-scalable but effective solution to eliminate misclassified messages. NotebookLM functioned as a general "outlier detector," identifying messages that deviated from their assigned topics. We uploaded our generated data, flagged irrelevant content, and iteratively removed problematic entries until achieving the desired quality. The final dataset contained 11,448 entries (Table 1). Future ELTEX implementations will integrate dedicated QA capabilities directly into the add-on for greater scalability and efficiency.
@@
4 Google Sheets Add-on
We specifically designed a Google Sheets add-on (using Apps Script [^3] https://developers.google.com/apps-script) to make the entire ELTEX pipeline accessible to non-technical users such as students or domain experts with minimal programming experience. This tool is already in use within our organization to create real-world datasets to meet internal needs.
Unlike Python-based packages — which require installing libraries, maintaining a local environment, and at least minimal scripting skills — or specialized online sandboxes (e.g., Hugging Face Spaces) that may feel unfamiliar to end-users, our add-on leverages a familiar and widely adopted platform: Google Sheets. This design choice reduces deployment overhead (no local installations needed), centralizes data management (Sheets inherently tracks revisions and supports collaboration), and offers a straightforward UI that many domain experts already use in their daily workflows.
Figure 2: Google Sheets Add-on Workflow. The logos of specific LLM providers are included here as examples only; any similarly capable model could be substituted for each of the depicted LLM services.
Figure 2 provides a high-level overview of the system. Below, we summarize the key steps:
  1. 1.
    User Parameters Input. The user opens a custom sidebar in Google Sheets and specifies the desired parameters for synthetic data generation: (a) topic (e.g., “cyberattacks”), (b) industry details (e.g., “blockchain”), (c) stakeholders (e.g., “exchanges”), (d) the dataset size, and (e) the generation temperature (defaulting to 0.8 based on our experiments, but fully adjustable by the user). Note: Because our pipeline may remove duplicates or near-duplicates in subsequent steps, the final retained sample count can be smaller than the user’s requested number.
    While some users provide seed data (e.g., a handful of real social media messages) for more precise context, others may enter textual descriptions of what they need, allowing the system to generate its initial examples.
    Additional Context (Optional). In addition to these parameters, users can supply textual fields labeled “Historical Events” or “General Knowledge” to guide indicator generation. When querying the LLMs for relevant indicators, the add-on automatically includes this optional background.
    Advanced Settings. By default, we select recommended up-to-date LLM providers and models (e.g., GPT-4o, Claude 3.5), but users can override these in the “Advanced Settings” panel, enabling them to choose or update providers/models according to their specific requirements.
  2. 2.
    Indicators Generation. Using the user parameters (and any optional historical context), the tool queries multiple LLMs (e.g., GPT-4o, Claude 3.5) to generate candidate domain indicators. We typically run these indicator-generation calls via a separate microservice to avoid Apps Script’s runtime limits. A final summarization step (using a frontier LLM) merges and refines these candidates into one concise list.
  3. 3.
    Combining Seed Data. If the user has provided seed data in the Sheet, the add-on includes them alongside the newly generated indicators and user-defined parameters to form the final prompt. If no seed data is available, the user’s textual descriptions and generated indicators alone guide the synthetic data generation.
  4. 4.
    Synthetic Data Generation. Rather than invoking LLM APIs directly from Apps Script — which has strict execution time limits — we dispatch requests to a separate backend (Cloudflare Workers Cloudflare (2025)) that handles the actual model calls. Our backend uses batch endpoints (e.g., OpenAI’s Batch API) to queue multiple prompt requests and process them together for efficiency and cost savings. This approach avoids the script timeout issue and reduces per-token costs. Once the batch results are ready, the backend returns the synthesized data to the add-on, which automatically appends them to a “Generated Data” sheet.
  5. 5.
    Deduplication. To remove near-duplicate data entries, the user can click “Deduplicate” in the Sheets menu. This triggers a Cloudflare Worker that runs the deduplication pipeline (Appendix E) and writes the cleaned data to a new sheet.
  6. 6.
    Data Export. Once satisfied, users can download the curated synthetic dataset (e.g., in CSV or JSON) directly from Google Sheets, integrating seamlessly with broader organizational workflows.
4.1 Scaling and System Bottlenecks
While our Google Sheets add-on offers a familiar interface and minimal deployment overhead, several technical constraints can arise at scale:
  • Apps Script Execution Time. Add-ons have a strict 30-second execution limit. To mitigate this, we offload heavy tasks (like LLM calls and deduplication) to external backends.
  • Cloudflare Workers Resource Limits. The backend runs on Cloudflare Workers, which also has finite CPU and memory quotas (e.g., a few hundred MB per request). To avoid exceeding these limits, we split massive generation jobs into multiple sequential batch requests, rather than submitting a single large payload.
Stateless vs. Stateful Workflows.
Each Cloudflare Worker invocation is stateless — no in-memory data persists once a request is completed. However, the system temporarily retains user-specific data: the add-on includes a user_id with each request, allowing our backend to associate multi-step operations with the correct spreadsheet session. Additionally, the deduplication service briefly stores embeddings keyed by user_id (for up to 24 hours) in a database. This limited retention helps streamline the pipeline while maintaining user privacy and conserving resources. Overall, the design combines the scalability of a stateless, serverless model with just enough short-term state to support multi-step workflows.
Cost Considerations.
As detailed in Appendix F, LLM usage fees dominate our total costs. However, we observe that OpenAI’s Batch API (or similar endpoints) can reduce per-token costs by up to 50% compared to synchronous calls. This batch processing approach offers higher rate limits but may delay completion by up to 24 hours, a potential bottleneck for users requiring immediate turnaround. In practice, most batches finish within a few minutes to a few hours, especially for moderate job sizes. Cloudflare Workers itself incurs only modest costs, and Apps Script imposes no additional fees beyond potential usage caps and timeouts.
5 Case Study
Model Performance Metrics Error Rates
Acc. Brier Recall F1 ROC False Pos. False Neg.
Gemma-2b 0.51 0.43 0.30 0.30 0.44 0.24 0.25
   w/ Real 0.65 0.31 0.47 0.61 0.73 0.03 0.32
   w/ ELTEX 0.77 0.16 0.78 0.76 0.85 0.13 0.11
   w/ Real + ELTEX 0.82 0.14 0.79 0.81 0.88 0.08 0.10
Llama-Primus 0.71 0.29 0.43 0.58 0.69 0.03 0.27
   w/ Merged 0.76 0.23 0.59 0.70 0.76 0.04 0.19
GPT-4o 0.84 0.10 0.71 0.81 0.94 0.02 0.14
Table 2: Performance on social media cyberattack detection. Best scores in bold, second-best underlined.
This section presents our evaluation of ELTEX-generated data for cybersecurity text classification, where each model classifies social media messages for potential relevance for blockchain attacks using a 0-1 scoring system. We focus on practical metrics relevant for industrial applications: classification performance, calibration quality using Brier score Brier (1950), and efficiency gains.
Table 2 summarizes the performance metrics for all evaluated models: the Gemma-2B base model, fine-tuned variants using LoRA Hu et al. (2022) (on real data, synthetic data and a hybrid approach) and GPT-4o as a reference point. We include cybersecurity-focused LLMs fine-tuned on the Primus dataset Yu et al. (2025) to contextualize our results.